OpenStudy (anonymous):

to make 5 sets of curtains which each set requires 2 1/3 yards how many of material are needed?

OpenStudy (anonymous):

5 x 2 1/3 turn 2 1/3 into an improper fraction 5/1 x 7/3 = 35/3 = 11 2/3 yards

OpenStudy (anonymous):

any questions ?